# GMIMS-DRAGONS: A Faraday Depth Survey of the Northern Sky Covering 350 to 1030 MHz

### v1

@@ 4.3 Base level Adjustments and Filtering
    We removed single-integration outliers to eliminate remaining RFI not flagged in the initial RFI masking (Section 4.1). These outliers were identified as data points having an absolute intensity value (in Stokes $I$, $Q$, or $U$) three times higher than the absolute intensity value of either of the two adjacent points in the time series, and were replaced by the mean of the adjacent data points.
@@
    We subtracted a daily median from each of Stokes $I$, $Q$ and $U$ at each frequency. This removed any overall frequency- and time-dependent instrumental offset. The calculations of the medians excluded a $5\arcdeg$ latitude range on either side of the Galactic midplane to avoid biasing the median values with strong Galactic emission. For Stokes $Q$ the daily median varied between $\sim$-200 and $\sim$400 Jy beam-1, and for Stokes $U$ between $\sim$-50 and $\sim$100 Jy beam-1 across the band, which corresponds to the Stokes $I$ to $Q$ and $U$ polarization leakage.
@@
    We subtracted a constant, fitted base level (average) from each scan in Stokes $I$, $Q$ and $U$ at each frequency to ensure a similar base level between adjacent scans and between East and West scans at the intersection points. For Stokes $Q$ and $U$, the large sinusoidal modulation by the parallactic angle along an East or West scan ensures that the data along each scan should average to zero. This is not the case for Stokes $I$, meaning that the zero-level is lost as a consequence of this step and a correction is required post-map-making. This step also removed the constant component of the ground contributions in Stokes $Q$ and $U$ (see Section 4.4).
@@
    We discarded scans with high rms fluctuations in either $Q$, $U$ or $I$ at each frequency. In the 350–600 MHz range, typically between 30 to 300 of each of the 2812 ${e}_{\text{high}}$ and 2802 ${e}_{\text{low}}$ observed scans9991406 ${e}_{\text{high}}$ and 1401 ${e}_{\text{low}}$ full $360\arcdeg$ scans yield 2812 ${e}_{\text{high}}$ and 2802 ${e}_{\text{low}}$ $180\arcdeg$ East and West scans. were discarded, while in the 600–1030 MHz range most channels had less than 30 scans discarded. Channels with more than $\sim$300 discarded scans at the low frequencies, and more than $\sim$100 discarded scans at the high frequencies were excluded from the final maps because of insufficient spatial sampling.
@@
4.4 Ground Correction
With the constant-elevation azimuth scan strategy used, each azimuth in an East or West scan (for each set of ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans) corresponds to a unique declination (see Figure 7). We begin with the assumption that the sky averages to zero in Stokes $Q$ and $U$ across a large RA range at each declination, and any remaining signal in a narrow declination bin is due to ground contamination from the corresponding azimuth. This is a reasonable assumption, given that we expect Stokes $Q$ and $U$ to fluctuate on scales much smaller than the full RA coverage. We averaged the data in $1\arcdeg$-wide declination bins separately for East, West, ${e}_{\text{high}}$ and ${e}_{\text{low}}$ sets of scans to produce a ground profile in Stokes $Q$ and $U$ at each frequency (examples in Figures 15 and 16). These profiles, which we subtracted from the data, represent the azimuth-dependent component of the ground contribution, while the constant component is already subtracted for each scan in Step 4 described in Section 4.3.
@@ 4.4 Ground Correction
The effectiveness of the ground correction improved when we calculated separate ground profiles for each of the three survey phases to account for seasonal variation, and subtracted these from the data accordingly. This is not unexpected, as ground properties at low frequencies vary rapidly with water content (Hall, 1979). The ground profiles for the three phases are shown as the red, yellow and blue lines in Figures 15 and 16. The dashed black line in each panel depicts the elevation profile of the horizon around the DRAO. Particularly in Stokes $Q$, there is a strong resemblance between the derived ground profiles and the physical features of the horizon. We expect the ground emission to be vertically polarized, therefore primarily appearing as positive Stokes $Q$. Since the constant component of the ground emission is already subtracted in Figures 15 and 16, these profiles do not show whether ground Stokes $Q$ is always positive. As well, on rough mountainous terrain the ground emission will be polarized perpendicular to the local ground surface, and there may be considerable Stokes $U$ and negative Stokes $Q$ components. These effects are more pronounced at lower frequencies.
Figure 17: Histograms showing the differences in Stokes $Q$ (Left) and Stokes $U$ (Right) between East and West high-elevation scans at their intersection points before (grey lines) and after (black lines) correcting for ground emission. The distributions generally become narrower and better centered on zero after correcting for the ground, showing the improved agreement between the East and West maps as a result of this step.
Since the East and West maps are observed in different ranges of azimuth, the ground contributions in the two maps are not the same. Therefore the East and West maps are not identical prior to ground correction, even though they contain the same sky signal. If our ground correction technique is effective, the difference between East and West maps after correction will be smaller than the difference before correction. Application of this test demonstrated that the differences were indeed reduced, but they were not reduced to zero, indicating that some ground emission remained in the maps. This proved to be a problem at later stages of processing (see Section 4.6). At selected frequencies, shown in Figure 17, we compared the distributions of the differences between the East and West values, with and without the ground correction applied, for each of Stokes $Q$ and $U$. The East-West difference distribution typically became 30% to 50% narrower with the ground correction applied, indicating substantially improved agreement between the East and West maps. However, at frequencies below $\sim 500$ MHz the uncertainties in the ground calculations still produce significant residual differences between the East and West maps. We incorporate these differences into error estimates for the maps as described in Section 5.3.
4.5 Ionospheric Faraday Rotation
@@
Figure 20: Histograms showing the differences in polarization angle, $\tau$, between East and West high-elevation scans (Left) and low-elevation scans (Right) at their intersection points before (grey lines) and after (black lines) correcting for ionospheric Faraday rotation. The percentages shown in each panel indicate the fraction of the data set that has less than $10\arcdeg$ difference between East and West, also highlighted by the blue shaded region on each histogram.
@@
As shown in Figure 19, despite being on the rising slope of the solar cycle, a significant fraction of the survey was observed with relatively low ionospheric RM. This is because the ionospheric Faraday rotation drops rapidly after sunset. For sets of scans conducted during the winter phase (P3; with more than 15 hours of observing possible on many nights), only the scans scheduled close to sunset and sunrise had ionospheric RM values of up to 3 rad m-2, while the majority of scans only suffered $\sim 0.5$ rad m-2 of ionospheric rotation. The highest ionospheric RMs that form spatially coherent patches in the maps occurred during the early summer phase, P1 (compare with Figure 8). In the low elevation scans, these patches coincide with the central regions of the Galaxy, where instrumental polarization is severe, and the data are in any case not trustworthy (see Section 5.3).
@@
The improvement to the maps provided by the ionospheric Faraday rotation correction can be quantified by comparing the polarization angles of the East and West sets of scans at their intersection points before and after derotating by the ionosphere. We illustrate this in Figure 20, which shows distributions of the differences between the East and West polarization angles, $\Delta\tau=\tau_{\text{E}}-\tau_{W}$ for a sample set of frequencies. For these plots, we excluded $\sim 20\%$ of the data using a minimum PI threshold, as low PI data are subject to increased noise in polarization angle. We also masked out the Galactic plane, which is dominated by instrumental polarization and is not expected to improve substantially from an ionospheric rotation correction. The resulting difference distributions generally become more peaked at $\tau=0\arcdeg$ after applying the correction. The fraction of the data points with less than $10\arcdeg$ difference between East and West polarization angles is indicated on each histogram. When the ionospheric correction is applied, this fraction increases by 3 to 6%, with the largest improvement occurring at the lowest frequencies (352 MHz in the examples plotted), as expected.
4.6 Basket-Weaving
@@
The survey was designed in the expectation that basket weaving would be employed to adjust scan base levels, and the S-PASS software package includes routines for that purpose. However, those routines were unsuccessful when applied to our data. We believe that this failure arose from our inability to remove ground emission completely from the data. If there is an error in the ground-emission estimate for a West scan characterized by ${\Delta}Q_{\text{ground}}$ and ${\Delta}U_{\text{ground}}$, that error vector is rotated by the ionospheric RM correction, and still remains an error. There may well be a similar error on the East scan, where both the ground emission and ionospheric RM are different. Such errors will vary from point to point along a scan. Basket-weaving calculates one number as a base-level correction for each scan and cannot compensate for errors that vary along a scan. The effect of these discrepancies is most pronounced at the lower frequencies, below $\sim$500 MHz. For a future data release, we plan to improve our estimate of ground emission by using the radiation pattern of the telescope calculated with an electromagnetic simulator to compute the ground contribution. For the current publication and data release, we quantify these errors and include them in uncertainty maps.
4.7 Destriping and Combining Maps
@@
In each frequency channel, for each of Stokes $Q$, $U$ and $I$, we binned the calibrated scans into $0.5\arcdeg\times 0.5\arcdeg$ pixels on an equatorial grid, separately for East and West scans as well as ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans, producing four independent maps. For each of ${e}_{\text{high}}$ and ${e}_{\text{low}}$, the East and West maps are duplicate maps of the same sky, and so their differences can be used to assess the processing and calibration steps.
@@
We adapted the Emerson & Graeve (1988) filtering algorithm to reduce any remaining scanning artifacts in each of the four maps. This algorithm was developed to work with two independent maps produced from sets of orthogonal straight-line scans. The 2D Fourier transform of each map has strong power in a line orthogonal to the scan direction, passing through the center of the transform (at low spatial frequencies). Filtering out power along the two orthogonal lines through the center removes the scanning artifacts. In the case of this survey, the East and West scans in equatorial coordinates are (i) not exactly orthogonal and (ii) deviate from straight lines over a significant declination range (see Figure 8). Thus we implemented a modification to the Emerson & Graeve (1988) algorithm, applying cone-shaped (instead of linear) filters in the Fourier transforms of the East and West maps. The edges of the cones account for the range of slopes encompassed by the curvature of the scan patterns and we tapered these edges using a Hanning window. For each of the ${e}_{\text{high}}$ and ${e}_{\text{low}}$ sets of maps, we combined East and West using weights derived from this filter. Finally, to produce a single map for each frequency and Stokes parameter, the ${e}_{\text{high}}$ and ${e}_{\text{low}}$ maps were combined using a linear weighting function applied in the overlapping declination range.
4.8 Maps Converted to Brightness Temperature
@@
Conversion to brightness temperature (K) was achieved by assuming a Gaussian main beam and calculating the main beam solid angle, $\Omega$, as a function of frequency, from the measured beamwidths (Figure 2), $\theta_{\text{HPBW}}$, as
@@
where $k$ is the Boltzmann constant and $\lambda$ is the observing wavelength. This conversion factor varies from $\sim$15 to $\sim$19 Jy beam-1 K-1, as shown in the bottom panel of Figure 6. The variation is due to the fluctuations in aperture efficiency across the band (Figure 6, top panel) and variations in beam efficiency (Figure 6, middle panel). The frequency-dependent conversion was applied to the maps produced in the final step of map-making. We publish the resulting Stokes $Q$, $U$ and $I$ frequency data cubes along with the Faraday depth cubes that we derive from them (Section 6). Examples of the spectra are shown in Figure 21 for three lines of sight, in which complex Faraday rotation effects are already evident in the Stokes $Q$ and $U$ fluctuations. Examples of Stokes $I$, $Q$, $U$, polarized intensity, and polarization angle maps at a single frequency (610 MHz) are shown in Figures 2226.
Figure 21: Frequency spectra for three sample lines of sight. The error bars shown on the Stokes $Q$ and $U$ values are from the error maps described in Section 5.3. The values shown here are from the data cubes convolved to a common resolution of $3.6\arcdeg$ for Faraday synthesis.
@@ 5 Data Quality Assessment
5.1 Total Intensity Assessment
@@
5.2 Polarization Comparison with Dwingeloo
This paper describes a survey of polarized radio emission over a large fraction of the sky, thus it is instructive to compare our survey with the previous best knowledge of the polarized radio sky in this frequency range. Brouw & Spoelstra (1976) published maps of the polarized emission at 408, 465, 610, 820, and 1411 MHz. The data were obtained with the Dwingeloo 25 m Telescope in the 1960s, and carefully calibrated in amplitude and corrected for instrumental polarization, for spurious polarized emission generated by radiation from the ground, and for Faraday rotation in the ionosphere. We refer to this dataset as the ‘Dwingeloo data’. The Dwingeloo data set was a good representation of the polarized sky in its frequency range, but its spatial sampling was very far from complete and its frequency sampling very sparse. Our work presents a survey that is fully Nyquist sampled on the sky, is fully sampled in frequency from 350 to 1030 MHz (to the extent that RFI at the DRAO permitted at the time the observations were made), and covers a much larger area of the sky. This makes mapping a range of scales of ISM structures in this frequency range now possible, in both the spatial and Faraday depth domains.
A comparison of the DRAGONS and Dwingeloo survey parameters is shown in Table 2 and a comparison between the two data sets is shown in Figure 28. To produce the comparison plots we calculated the mean values of DRAGONS Stokes $Q$ and $U$ within the DRAGONS beam at the location of each Dwingeloo pointing. We then calculated PI and polarization angle from these mean $Q$ and $U$ values. The polarization angles are ‘un-wrapped’ by constraining the differences between the DRAGONS and Dwingeloo angles to be less than $90\arcdeg$ and adjusting the angles by $180\arcdeg$ accordingly. We fitted a linear equation to each scatter plot, using data points for which DRAGONS PI is above three times the noise level, and calculated the Pearson correlation coefficient, $R$. There is overall good correlation between DRAGONS and Dwingeloo both in Stokes $Q$ and $U$, as well as in PI ($R$ values of 0.6 to 0.9). The polarization angle agreement is even stronger ($R$ mostly above 0.9). The poorer correlations at 408 and 465 MHz compared to 610 and 820 MHz are likely in part caused by the uncertainties in the DRAGONS maps, which also tend to be higher at the lower frequencies (see Section 5.3).

### v2

@@ IV.3 Base level Adjustments and Filtering
    We removed single-integration outliers to eliminate remaining RFI not flagged in the initial RFI masking (Section IV.1). These outliers were identified as data points having an absolute intensity value (in Stokes $I$, $Q$, or $U$) 3 times higher than the absolute intensity value of either of the two adjacent points in the time series, and were replaced by the mean of the adjacent data points.
@@
    We subtracted a daily median from each of Stokes $I$, $Q$ and $U$ at each frequency. This removed any overall frequency- and time-dependent instrumental offset. The calculations of the medians excluded a $5\arcdeg$ latitude range on either side of the Galactic midplane to avoid biasing the median values with strong Galactic emission. For Stokes $Q$ the daily median varied between $\sim$-200 and $\sim$400 Jy beam-1, and for Stokes $U$ the daily median varied between $\sim$-50 and $\sim$100 Jy beam-1 across the band, which corresponds to the Stokes $I$ to $Q$ and $U$ polarization leakage.
@@
    We subtracted a constant, fitted base level (average) from each scan in Stokes $I$, $Q$ and $U$ at each frequency to ensure a similar base level between adjacent scans and between east and west scans at the intersection points. For Stokes $Q$ and $U$, the large sinusoidal modulation by the parallactic angle along an east or west scan ensures that the data along each scan should average to zero. This is not the case for Stokes $I$, meaning that the zero-level is lost as a consequence of this step and a correction is required after map-making. This step also removed the constant component of the ground contributions in Stokes $Q$ and $U$ (see Section IV.4).
@@
    We discarded scans with high rms fluctuations in either $Q$, $U$ or $I$ at each frequency. In the 350–600 MHz range, typically between 30 and 300 of each of the 2812 ${e}_{\text{high}}$ and 2802 ${e}_{\text{low}}$ observed scans9991406 ${e}_{\text{high}}$ and 1401 ${e}_{\text{low}}$ full $360\arcdeg$ scans yield 2812 ${e}_{\text{high}}$ and 2802 ${e}_{\text{low}}$ $180\arcdeg$ east and west scans. were discarded, while in the 600–1030 MHz range most channels had less than 30 scans discarded. Channels with more than $\sim$300 discarded scans at the low frequencies, and more than $\sim$100 discarded scans at the high frequencies were excluded from the final maps because of insufficient spatial sampling.
IV.4 Ground Correction
With the constant-elevation azimuth scan strategy used, each azimuth in an east or west scan (for each set of ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans) corresponds to a unique declination (see Figure 7). We begin with the assumption that the sky averages to zero in Stokes $Q$ and $U$ across a large RA range at each declination, and any remaining signal in a narrow declination bin is due to ground contamination from the corresponding azimuth. This is a reasonable assumption, given that we expect Stokes $Q$ and $U$ to fluctuate on scales much smaller than the full RA coverage. We averaged the data in $1\arcdeg$-wide declination bins separately for east, west, ${e}_{\text{high}}$ and ${e}_{\text{low}}$ sets of scans to produce a ground profile in Stokes $Q$ and $U$ at each frequency (examples in Figures 15 and 16). These profiles, which we subtracted from the data, represent the azimuth-dependent component of the ground contribution, while the constant component is already subtracted for each scan in Step 4 described in Section IV.3.
The effectiveness of the ground correction improved when we calculated separate ground profiles for each of the three survey phases to account for seasonal variation, and subtracted these from the data accordingly. This is not unexpected, as ground properties at low frequencies vary rapidly with water content (Hall, 1979). The ground profiles for the three phases are shown as the red, yellow, and blue lines in Figures 15 and 16. The dashed black line in each panel depicts the elevation profile of the horizon around the DRAO. Particularly in Stokes $Q$, there is a strong resemblance between the derived ground profiles and the physical features of the horizon.
@@ IV.4 Ground Correction
We expect the ground emission to be vertically polarized, therefore primarily appearing as positive Stokes $Q$. Since the constant component of the ground emission is already subtracted in Figures 15 and 16, these profiles do not show whether ground Stokes $Q$ is always positive. As well, on rough mountainous terrain the ground emission will be polarized perpendicular to the local ground surface, and there may be considerable Stokes $U$ and negative Stokes $Q$ components. These effects are more pronounced at lower frequencies.
Since the east and west maps are observed in different ranges of azimuth, the ground contributions in the two maps are not the same. Therefore the east and west maps are not identical prior to ground correction, even though they contain the same sky signal. If our ground correction technique is effective, the difference between east and west maps after correction will be smaller than the difference before correction. Application of this test demonstrated that the differences were indeed reduced, but they were not reduced to zero, indicating that some ground emission remained in the maps. This proved to be a problem at later stages of processing (see Section IV.6). At selected frequencies, shown in Figure 17, we compared the distributions of the differences between the east and west values, with and without the ground correction applied, for each of Stokes $Q$ and $U$. The east-west difference distribution typically became 30% to 50% narrower with the ground correction applied, indicating substantially improved agreement between the east and west maps. However, at frequencies below $\sim 500$ MHz the uncertainties in the ground calculations still produce significant residual differences between the east and west maps. We incorporate these differences into error estimates for the maps as described in Section V.3.
Figure 17: Histograms showing the differences in Stokes $Q$ (Left) and Stokes $U$ (Right) between east and west high-elevation scans at their intersection points before (gray lines) and after (black lines) correcting for ground emission. The distributions generally become narrower and better centered on zero after correcting for the ground, showing the improved agreement between the east and west maps as a result of this step.
IV.5 Ionospheric Faraday Rotation
@@
As shown in Figure 19, despite being on the rising slope of the solar cycle, a significant fraction of the survey was observed with relatively low ionospheric RM. This is because the ionospheric Faraday rotation drops rapidly after sunset. For sets of scans conducted during the winter phase (P3; with more than 15 hours of observing possible on many nights), only the scans scheduled close to sunset and sunrise had ionospheric RM values of up to 3 rad m-2, while the majority of scans only suffered $\sim 0.5$ rad m-2 of ionospheric rotation. The highest ionospheric RMs that form spatially coherent patches in the maps occurred during the early summer phase, P1 (compare with Figure 8). In the low elevation scans, these patches coincide with the central regions of the Galaxy, where instrumental polarization is severe, and the data are in any case not trustworthy (see Section V.3).
Figure 20: Histograms showing the differences in polarization angle, $\tau$, between east and west high-elevation scans (Left) and low-elevation scans (Right) at their intersection points before (gray lines) and after (black lines) correcting for ionospheric Faraday rotation. The percentages shown in each panel indicate the fraction of the data set that has less than $10\arcdeg$ difference between east and west, also highlighted by the blue shaded region on each histogram.
@@
The improvement to the maps provided by the ionospheric Faraday rotation correction can be quantified by comparing the polarization angles of the east and west sets of scans at their intersection points before and after derotating by the ionosphere. We illustrate this in Figure 20, which shows distributions of the differences between the east and west polarization angles, $\Delta\tau=\tau_{\text{E}}-\tau_{W}$ for a sample set of frequencies. For these plots, we excluded $\sim 20\%$ of the data using a minimum PI threshold, as low PI data are subject to increased noise in polarization angle. We also masked out the Galactic plane (all data within $|b|<5\arcdeg$), which is dominated by instrumental polarization and is not expected to improve substantially from an ionospheric rotation correction. The resulting difference distributions generally become more peaked at $\tau=0\arcdeg$ after applying the correction. The fraction of the data points with less than $10\arcdeg$ difference between east and west polarization angles is indicated on each histogram. When the ionospheric correction is applied, this fraction increases by 3 to 6%, with the largest improvement occurring at the lowest frequencies (352 MHz in the examples plotted), as expected.
IV.6 Basket-Weaving
@@
The survey was designed in the expectation that basket weaving would be employed to adjust scan base levels, and the S-PASS software package includes routines for that purpose. However, those routines were unsuccessful when applied to our data. We believe that this failure arose from our inability to remove ground emission completely from the data. If there is an error in the ground emission estimate for a west scan characterized by ${\Delta}Q_{\text{ground}}$ and ${\Delta}U_{\text{ground}}$, that error vector is rotated by the ionospheric RM correction, and still remains an error. There may well be a similar error on the east scan, where both the ground emission and ionospheric RM are different. Such errors will vary from point to point along a scan. Basket-weaving calculates one number as a base-level correction for each scan and cannot compensate for errors that vary along a scan. The effect of these discrepancies is most pronounced at the lower frequencies, below $\sim$500 MHz. For a future data release, we plan to improve our estimate of ground emission by using the radiation pattern of the telescope calculated with an electromagnetic simulator to compute the ground contribution. For the current publication and data release, we quantify these errors and include them in uncertainty maps.
IV.7 Destriping and Combining Maps
@@
In each frequency channel, for each of Stokes $Q$, $U$, and $I$, we binned the calibrated scans into $0.5\arcdeg\times 0.5\arcdeg$ pixels on an equatorial grid, separately for east and west scans as well as ${e}_{\text{high}}$ and ${e}_{\text{low}}$ scans, producing four independent maps. For each of ${e}_{\text{high}}$ and ${e}_{\text{low}}$, the east and west maps are duplicate maps of the same sky, and so their differences can be used to assess the processing and calibration steps.
@@
We adapted the Emerson and Graeve (1988) filtering algorithm to reduce any remaining scanning artifacts in each of the four maps. This algorithm was developed to work with two independent maps produced from sets of orthogonal straight-line scans. The 2D Fourier transform of each map has strong power in a line orthogonal to the scan direction, passing through the center of the transform (at low spatial frequencies). Filtering out power along the two orthogonal lines through the center removes the scanning artifacts. In the case of this survey, the east and west scans in equatorial coordinates are (i) not exactly orthogonal and (ii) deviate from straight lines over a significant declination range (see Figure 8). Thus we implemented a modification to the Emerson and Graeve (1988) algorithm, applying cone-shaped (instead of linear) filters in the Fourier transforms of the east and west maps. The edges of the cones account for the range of slopes encompassed by the curvature of the scan patterns and we tapered these edges using a Hanning window. For each of the ${e}_{\text{high}}$ and ${e}_{\text{low}}$ sets of maps, we combined east and west using weights derived from this filter. Finally, to produce a single map for each frequency and Stokes parameter, the ${e}_{\text{high}}$ and ${e}_{\text{low}}$ maps were combined using a linear weighting function applied in the overlapping declination range.
IV.8 Maps Converted to Brightness Temperature
@@
The conversion to brightness temperature (K) was achieved by assuming a Gaussian main beam and calculating the main beam solid angle, $\Omega$, as a function of frequency, from the measured beamwidths (Figure 2), $\theta_{\text{HPBW}}$, as
@@
where $k$ is the Boltzmann constant and $\lambda$ is the observing wavelength. This conversion factor varies from $\sim$15 to $\sim$19 Jy beam-1 K-1, as shown in the bottom panel of Figure 6. The variation is due to the fluctuations in aperture efficiency across the band (Figure 6, top panel) and variations in beam efficiency (Figure 6, middle panel). The frequency-dependent conversion was applied to the maps produced in the final step of map-making. We publish the resulting Stokes $Q$, $U$ and $I$ frequency data cubes along with the Faraday depth cubes that we derive from them (Section VI). Examples of the spectra are shown in Figure 21 for three lines of sight, in which complex Faraday rotation effects are already evident in the Stokes $Q$ and $U$ fluctuations. Examples of Stokes $I$, $Q$, $U$, PI, and polarization angle maps at a single frequency (610 MHz) are shown in Figures 2226.
V Data Quality Assessment
Figure 21: Frequency spectra for three sample lines of sight. The error bars shown on the Stokes $Q$ and $U$ values are from the error maps described in Section V.3. The values shown here are from the data cubes convolved to a common resolution of $3.6\arcdeg$ for Faraday synthesis.
@@ V Data Quality Assessment
V.1 Total Intensity Assessment
@@
V.2 Polarization Comparison with Dwingeloo
This paper describes a survey of polarized radio emission over a large fraction of the sky, thus it is instructive to compare our survey with the previous best knowledge of the polarized radio sky in this frequency range. Brouw and Spoelstra (1976) published maps of the polarized emission at 408, 465, 610, 820, and 1411 MHz. The data were obtained with the Dwingeloo 25 m Telescope in the 1960s, and carefully calibrated in amplitude and corrected for instrumental polarization, for spurious polarized emission generated by radiation from the ground, and for Faraday rotation in the ionosphere. We refer to this dataset as the ‘Dwingeloo data’. The Dwingeloo data set was a good representation of the polarized sky in its frequency range, but its spatial sampling was very far from complete and its frequency sampling very sparse. Our work presents a survey that is fully Nyquist sampled on the sky, is fully sampled in frequency from 350 to 1030 MHz (to the extent permitted by RFI at the DRAO at the time the observations were made), and covers a much larger area of the sky. This makes mapping a range of scales of ISM structures in this frequency range now possible, in both the spatial and Faraday depth domains.
A comparison of the DRAGONS and Dwingeloo survey parameters is shown in Table 2 and a comparison between the two data sets is shown in Figure 28. To produce the comparison plots we calculated the mean values of DRAGONS Stokes $Q$ and $U$ within the DRAGONS beam at the location of each Dwingeloo pointing. We then calculated PI and polarization angle from these mean $Q$ and $U$ values. The polarization angles are ‘un-wrapped’ by constraining the differences between the DRAGONS and Dwingeloo angles to be less than $90\arcdeg$ and adjusting the angles by $180\arcdeg$ accordingly. We fitted a linear equation to each scatter plot, using data points for which DRAGONS PI is above 3 times the noise level, and calculated the Pearson correlation coefficient, $R$. There is overall good correlation between DRAGONS and Dwingeloo both in Stokes $Q$ and $U$, as well as in PI ($R$ values of 0.6 to 0.9). The polarization angle agreement is even stronger ($R$ mostly above 0.9). The poorer correlations at 408 and 465 MHz compared to 610 and 820 MHz are likely in part caused by the uncertainties in the DRAGONS maps, which also tend to be higher at the lower frequencies (see Section V.3).